OpenStudy (anonymous):

if the accepted value of pi is 3.1416, what is the percent error of the experimental value 3.15?

OpenStudy (aum):

Percentage error = |(Experimental value - Theoretical/Accepted value)| / (Theoretical/Accepted value) * 100

OpenStudy (aum):

Percentage error = \(\Large \frac {|3.15 - 3.1416|}{3.1416} \times 100 = ?\)

OpenStudy (anonymous):

So my answer would be.. 0.26738?